# FlashKAT: Understanding and Addressing Performance Bottlenecks in the Kolmogorov-Arnold Transformer

### v1

@@ 3 Beyond FLOPs: Understanding Performance Bottlenecks of KAT
Focusing on the memory side for backward passes, we note that Table 2 reports very low L1, L2, and HBM throughput. At first glance, this is counterintuitive because a memory-bound process would usually mean a saturated memory bandwidth. However, further reasoning about the underlying hardware can provide insights. A GPU consists of a number of streaming multiprocessors (SMs). Typically, 32 threads are grouped in a warp and executed synchronously. Multiple warps can be scheduled for execution on an SM in a context-switching fashion. When one warp is waiting for a memory transfer, the SM can switch to doing computation for a different "ready" warp (i.e., not waiting for memory), thus hiding the memory latency. However, suppose there are not enough ready warps. Then, the warp scheduler cannot entirely hide long latency or high memory contention, like with atomic operations that create sequential memory accesses. If an SM does not make adequate use of the memory hierarchy consisting of registers, shared memory, and global memory (a mapped memory composed of L1 cache, L2 cache, and HBM), the latency for each transfer becomes even more severe. For reference, on an A100, shared memory, L1, L2, and HBM have a latency of 29.0, 37.9, 261.5, and 466.3 clock cycles, respectively luo2024benchmarking . Assuming that the group-wise rational fails to hide latency, it makes sense that we would witness low memory utilization but still have a memory bottleneck. To validate this assumption, we examine the warp state statistics for the group-wise rational backward pass using Nvidia Nsight Compute with the same configuration as before. These statistics provide the average number of cycles each warp stays in a specific state for each instruction the kernel executes. We report the warp state statistics in Figure 2.
@@
Figure 2 demonstrates that each warp spends a majority of its time waiting for various memory stalls to conclude. Each stall reported is related to a memory transfer, excluding the "Selected" state, where the warp is selected to issue an instruction (i.e., computing). For instance, the "Stall Long Scoreboard", a stall on a memory transfer from global memory, accounts for 412x more time than the time spent in the "Selected" state. As such, it becomes clear that the group-wise rational slowdown is primarily due to memory stalls from memory transfers. This calls for a new method that can be aware of the memory hierarchy and avoid memory operations that create contention and restrict the latency hiding capabilities of the warp scheduler.
@@ 4 FlashKAT
In the previous section, we identified that the KAT has a memory bottleneck in the backward pass of the group-wise rational. In this section, we first outline the computations required for this operation, then provide an analysis of the standard group-wise rational used in the state-of-the-art KAT to pin down the source of the bottleneck, and finally present FlashKAT, our proposed Kolmogorov-Arnold Transformer that builds on an optimized group-wise rational that efficiently leverages gradient accumulation and memory hierarchy to minimize global memory accesses.
4.1 Gradient Computations
@@
Since group-wise rational, expressed in Equation 6 as $F(\cdot)$, is an element-wise operation composed of learnable parameters $a_{i}$ and $b_{j}$ for the numerator and the denominator, the backward pass requires us to compute the gradients $\frac{\partial F}{\partial a_{i}}$, $\frac{\partial F}{\partial b_{j}}$, and $\frac{\partial F}{\partial x}$. Suppose that $A(x)=b_{1}x+...,+b_{n}x^{n}$, $\frac{\partial P(x)}{\partial x}=a_{1}+2a_{2}x+...+ma_{m}x^{m-1}$, and $\frac{\partial Q(x)}{\partial x}=\frac{\partial(1+|A(x)|)}{\partial A(x)}\cdot\frac{\partial A(x)}{\partial x}=\frac{A(x)}{|A(x)|}(b_{1}+2b_{2}x+...+nb_{n}x^{n-1})$ then the gradients for each is provided in Equation 7:
$\frac{\partial F}{\partial a_{i}}=\frac{x^{i}}{Q(x)},\quad\frac{\partial F}{\partial b_{j}}=-x^{j}\frac{A(x)}{|A(x)|}\frac{P(x)}{Q(x)^{2}},\quad\frac{\partial F}{\partial x}=\frac{\partial P(x)}{\partial x}\frac{1}{Q(x)}-\frac{\partial Q(x)}{\partial x}\frac{P(x)}{Q^{2}(x)}$ (7)
@@
Equation 7 provides the element-wise gradients; however, since each group of size $d_{g}$, each batch of size $B$ and each element in the sequence of size $N$ share an identical set of coefficients, the contributions of each must be accumulated on $\frac{\partial F}{\partial a_{i}}$ and $\frac{\partial F}{\partial b_{j}}$. Let us denote $\frac{\partial F}{\partial a_{g,i}}$ and $\frac{\partial F}{\partial b_{g,j}}$ as the gradient for coefficient $i$ and $j$ of group $g$. Then we can express this accumulation in Equation 8:
$\frac{\partial F}{\partial a_{g,i}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial F(X_{i^{\prime},j^{\prime},g\cdot d_{g}+k^{\prime}})}{\partial a_{g,i}}\text{, and }\frac{\partial F}{\partial b_{g,j}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial F(X_{i^{\prime},j^{\prime},g\cdot d_{g}+k^{\prime}})}{\partial b_{g.j}}$ (8)
@@
There are multiple different methods for accumulating these gradients in a kernel, and identifying the optimal accumulation procedure is crucial for a fast training algorithm. We will now demonstrate where the standard group-wise rational method falls short in this regard.
4.2 Analysis of KAT group-wise rational
@@
Provided an input $\mathbf{X}\in\mathbb{R}^{B\times N\times d}$, an output gradient $\mathbf{dO}\in\mathbb{R}^{B\times N\times d}$ and coefficients $\mathbf{A}\in\mathbb{R}^{n_{g}\times m}$ and $\mathbf{B}\in\mathbb{R}^{n_{g}\times n}$, we want to compute the element-wise gradients $\frac{\partial F}{\partial\mathbf{A}}$ ($\mathbf{dA}$), $\frac{\partial F}{\partial\mathbf{B}}$ ($\mathbf{dB})$, and $\frac{\partial F}{\partial\mathbf{X}}$ ($\mathbf{dX}$). In Algorithm 1, we provide the standard method for computing each of these gradients in the KAT. The algorithm is structured sequentially, and all the on-chip computation is done in the subroutine calls, following Equation 7, for readability.
Algorithm 1 The KAT group-wise rational backward pass.
1:$\mathbf{X},\mathbf{dO}\in\mathbb{R}^{B\times N\times d}$,  $\mathbf{A}\in\mathbb{R}^{n_{g}\times m}$,   $\mathbf{B}\in\mathbb{R}^{n_{g}\times n}$ in global memory.
@@
3:Divide $\mathbf{X}$ and dO into $T=\lceil(B\cdot N\cdot d/S_{block}\rceil$ blocks $\mathbf{X_{1}},...,\mathbf{X_{T}}$ and $\mathbf{dO_{1}},...,\mathbf{dO_{T}}$ each of size $S_{\text{block}}$.
@@
10:        Load $A_{k\cdot m,1},...,A_{k\cdot m,m}$ and $B_{k\cdot n,1},..,B_{k\cdot n,n}$ from global memory
11:        $dX_{i,j}\leftarrow\textbf{Compute\_dX}(A_{k\cdot m,1},...,A_{k\cdot m,m},B_{k\cdot n,1},..,B_{k\cdot n,n},X_{i,j},dO_{i,j})$
12:        $\mathbf{dA_{k\cdot m}}^{\prime}\leftarrow\textbf{Compute\_dA}(A_{k\cdot m,1},...,A_{k\cdot m,m},B_{k\cdot n,1},..,B_{k\cdot n,n},X_{i,j},dO_{i,j})$
13:        $\mathbf{dB_{k\cdot n}}^{\prime}\leftarrow\textbf{Compute\_dB}(A_{k\cdot m,1},...,A_{k\cdot m,m},B_{k\cdot n,1},..,B_{k\cdot n,n},X_{i,j},dO_{i,j})$
14:        Load $\mathbf{dA_{k\cdot m}}$ from global memory, set $\mathbf{dA_{k\cdot m}}\leftarrow\mathbf{dA_{k\cdot m}}+\mathbf{dA^{\prime}_{k\cdot m}}$, write back to global memory
15:        Load $\mathbf{dB_{k\cdot n}}$ from global memory, set $\mathbf{dB_{k\cdot n}}\leftarrow\mathbf{dB_{k\cdot n}}+\mathbf{dB^{\prime}_{k\cdot n}}$, write back to global memory
@@
At a high level, Algorithm 1 will iterate over the $T$ blocks in the grid, and for each block there will be $S_{block}$ threads that will compute the associated gradients for $\mathbf{dA}$, $\mathbf{dB}$, $\mathbf{dX}$. In the case of $\mathbf{dA}$ and $\mathbf{dB}$, for each of these computations, the thread will perform an atomic add, which requires reading from global memory, accumulating the value, and then storing the result back to global memory as an indivisible unit of work. However, since each of these threads needs to perform an atomic add for each of the $m+n$ coefficients, the atomic adds must occur sequentially. As such, for each of these atomic adds, a stall occurs. Since multiple threads within the same block and across different blocks may be writing to the same location, there will be extreme resource contention. Such resource contention inevitably leads to high amounts of latency and stalls without utilizing the available bandwidth, as we demonstrated in Section 1.
@@
Let us analyze the total number of global memory accesses incurred by Algorithm 1. We can see that each $X_{i}$ and $dO_{i}$ requires $S_{block}$ reads from global memory. Then writing $dX_{i}$ back to global memory will require $S_{block}$ writes to global memory. Since these 3 memory actions occur $T$ times they will require $3TS_{\text{block}}=3\frac{B\cdot N\cdot d}{S_{\text{block}}}\cdot S_{\text{block}}=3\cdot B\cdot N\cdot d$ global memory accesses. Regarding the coefficients, each element in the output $dX$ requires loading $m+n$ coefficients from global memory. It also requires the aforementioned atomic add, which includes an additional $(m+n)$ reads and $(m+n)$ writes to global memory. Thus, since there are $T$ blocks of $S_{\text{block}}$ threads, there will be $TS_{\text{block}}\cdot 3(m+n)=3(m+n)\cdot B\cdot N\cdot d$ accesses. As a result, there are $3(m+n+1)\cdot B\cdot N\cdot d$ global memory accesses in total. To re-emphasize, the memory accesses for the coefficients for the atomic add are much worse due to the additional contention with other threads.
4.3 FlashKAT group-wise rational
@@
We now present an optimized method for computing the backward pass of the group-wise rational in Algorithm 2. To reduce global memory accesses, we restructure the grid from 1D to 2D. Such a restructuring mirrors the format of Equation 8, allowing an easier mapping of it to the algorithm. The first dimension of our reconstructed grid is allocated for resolving the $B$ and $N$ summations, whereas the second dimension is responsible for the $d_{g}$ summation. The size of this first dimension in the grid is $B\cdot N/S_{\text{block}}$, whereas the size of the second dimension is $n_{g}$ (the total number of groups in GR-KAN).
@@
Algorithm 2 The FlashKAT group-wise rational backward pass.
1:$\mathbf{X},\mathbf{dO}\in\mathbb{R}^{B\times N\times d}$,  $\mathbf{A}\in\mathbb{R}^{n_{g}\times m}$,   $\mathbf{B}\in\mathbb{R}^{n_{g}\times n}$ in global memory.
2:Set block sizes to $S_{block}$ and $d_{g}=d/n_{g}$.
@@
3:Divide $\mathbf{X}$ and $\mathbf{dO}$ into $\lceil(B\cdot N\cdot d/(S_{block}\cdot d_{g})\rceil$ blocks $\mathbf{X_{i,j}}\in\mathbb{R}^{S_{\text{block}}\times d_{g}}$ and $\mathbf{dO_{i,j}}\in\mathbb{R}^{S_{\text{block}}\times d_{g}}$ of a $T\times n_{g}$ grid where $T=\lceil(B\cdot N/S_{block}\rceil$.
@@
9:        Load $\mathbf{X_{i.j}}$ from global memory
@@
14:               $dX_{i,j,k,l}\leftarrow\textbf{Compute\_dX}(A_{j,1},...,A_{j,m},B_{j,1},..,B_{j,n},X_{i,j,k,l},dO_{i,j,k,l})$
15:               $\mathbf{dA_{j}}^{\prime}\leftarrow\mathbf{dA_{j}}^{\prime}+\textbf{Compute\_dA}(A_{j,1},...,A_{j,m},B_{j,1},..,B_{j,n},X_{i,k},dO_{i,j,k,l})$
16:               $\mathbf{dB_{j}}^{\prime}\leftarrow\mathbf{dB_{j}}^{\prime}+\textbf{Compute\_dB}(A_{j,1},...,A_{j,m},B_{j,1},..,B_{j,n},X_{i,j,k,l},dO_{i,j,k,l})$
@@
19:        Load $\mathbf{dA_{j}}$ from global memory, set $\mathbf{dA_{j}}\leftarrow\mathbf{dA_{j}}+\mathbf{dA_{j}}^{\prime}$, write back to global memory
20:        Load $\mathbf{dB_{j}}$ from global memory, set $\mathbf{dB_{j}}\leftarrow\mathbf{dB_{j}}+\mathbf{dB_{j}}^{\prime}$, write back to global memory
@@
Algorithm 2 does not change the memory accesses for reads and writes for $\mathbf{dX},\mathbf{X}$ and $\mathbf{dO}$; however, it heavily influences them for $\mathbf{A}$, $\mathbf{B}$, $\mathbf{dA}$, and $\mathbf{dB}$. At a high level, we can see that our algorithm iterates over the grid of $T\cdot n_{g}$ blocks. For each of these blocks, we load a single group of coefficients for $\mathbf{A_{j}}$ and $\mathbf{B_{j}}$. Then, we use this single set of coefficients for all computations in the block. As a result, rather than each thread performing an atomic add for each value in $\mathbf{dA_{j}}$ and $\mathbf{dB_{j}}$, the algorithm can instead accumulate all the contributions for $\mathbf{dA_{j}}$ and $\mathbf{dB_{j}}$ in the block before doing an atomic add. As a result, the total amount of memory contention and global memory accesses reduce.
We will now analyze the total global memory accesses for Algorithm 2. A block will first load one set of coefficients from the global memory associated with the group to which the block is assigned. Since there is only 1 set of coefficients per block, this action will only require $m+n$ accesses. Then the algorithm will load the associated $\mathbf{X_{i,j}}$ and $\mathbf{dO_{i,j}}$ from global memory, each of which is of size $S_{\text{block}}\cdot d_{g}$. After computing $\mathbf{dX_{i,j}}$, its $S_{\text{block}}\cdot d_{g}$ values will be written back to global memory. Then, an atomic add will load $m+n$ values from $\mathbf{dA_{j}}$ and $\mathbf{dB_{j}}$, accumulate them with $\mathbf{dA_{j}}^{\prime}$ and $\mathbf{dB_{j}}^{\prime}$, and write them back to global memory. Since all these operations occur across the $T\cdot n_{g}$ blocks, there will be $T\cdot n_{g}(3S_{\text{block}}\cdot d_{g}+3(m+n))=3\frac{B\cdot N}{S_{\text{block}}}\cdot\frac{d}{d_{g}}(S_{\text{block}}\cdot d_{g}+(m+n))=3(\frac{m+n}{S_{\text{block}}\cdot d_{g}}+1)\cdot B\cdot N\cdot d$ global memory access. Compared to Algorithm 1, we can see that Algorithm 2 reduces the number of global memory accesses and subsequently the number of atomic adds by a factor of $\frac{1}{S_{\text{block}}\cdot d_{g}}$.

### v2

@@ 3 Beyond FLOPs: Understanding Performance Bottlenecks of KAT
Focusing on the memory side for backward passes, we note that Table 2 reports very low L1, L2, and HBM throughput. At first glance, this is counterintuitive because a memory-bound process would usually mean a saturated memory bandwidth. However, further reasoning about the underlying hardware can provide insights. A GPU consists of a number of streaming multiprocessors (SMs). Typically, 32 threads are grouped in a warp and executed synchronously. Multiple warps can be scheduled for execution on an SM in a context-switching fashion. When one warp is waiting for a memory transfer, the SM can switch to doing computation for a different ”ready” warp (i.e., not waiting for memory), thus hiding the memory latency. However, suppose there are not enough ready warps. Then, the warp scheduler cannot entirely hide long latency or high memory contention, like with atomic operations that create sequential memory accesses. If an SM does not make adequate use of the memory hierarchy consisting of registers, shared memory, and global memory (a mapped memory composed of L1 cache, L2 cache, and HBM), the latency for each transfer becomes even more severe. For reference, on an A100, shared memory, L1, L2, and HBM have a latency of 29.0, 37.9, 261.5, and 466.3 clock cycles, respectively (luo2024benchmarking). Assuming that the group-wise rational function evaluation fails to hide latency, it makes sense that we would witness low memory utilization but still have a memory bottleneck. To validate this assumption, we examine the warp state statistics for the group-wise rational function backward pass using Nvidia Nsight Compute with the same configuration as before. These statistics provide the average cycles each warp stays in a given state for each instruction the kernel executes. We report the warp state statistics in Figure 2.
@@
Figure 2 demonstrates that each warp spends the majority of its time waiting for various memory stalls to conclude. Each stall reported is related to a memory transfer, excluding the ”Selected” state, where the warp is selected to issue an instruction (i.e., computing). For instance, the ”Stall Long Scoreboard”, a stall on a memory transfer from global memory, accounts for 412x more time than the time spent in the ”Selected” state. As such, it becomes clear that the group-wise rational function slowdown is primarily due to memory stalls from memory transfers. This calls for a new method that can be aware of the memory hierarchy and avoid memory operations that create contention and restrict the latency hiding capabilities of the warp scheduler.
@@ 4 FlashKAT
In the previous section, we identified that the KAT has a memory bottleneck in the backward pass of the group-wise rational function, a discovery that previous approaches had not recognized (yang2024kolmogorov; molina2019pad; delfosse2021recurrent). In this section, we first outline the computations required for this operation, then provide an analysis of the standard group-wise rational function used in the state-of-the-art KAT to pin down the source of the bottleneck, and finally present FlashKAT, our proposed Kolmogorov-Arnold Transformer that builds on an optimized group-wise rational function that efficiently leverages gradient accumulation and memory hierarchy to minimize global memory accesses.
Gradient Computations
@@
Since group-wise rational function, expressed in Equation 6 as $F(\cdot)$, is an element-wise operation composed of learnable parameters $a_{i}$ and $b_{j}$ for the numerator and the denominator, the backward pass requires us to compute the gradients $\frac{\partial F}{\partial a_{i}}$, $\frac{\partial F}{\partial b_{j}}$, and $\frac{\partial F}{\partial x}$. Suppose that $A(x)=b_{1}x+...,+b_{n}x^{n}$, $\frac{\partial P(x)}{\partial x}=a_{1}+2a_{2}x+...+ma_{m}x^{m-1}$, and $\frac{\partial Q(x)}{\partial x}=\frac{\partial(1+|A(x)|)}{\partial A(x)}\cdot\frac{\partial A(x)}{\partial x}=\frac{A(x)}{|A(x)|}(b_{1}+2b_{2}x+...+nb_{n}x^{n-1})$ then the gradients for each are
$\frac{\partial F}{\partial a_{i}}=\frac{x^{i}}{Q(x)},$ (7)
$\frac{\partial F}{\partial b_{j}}=-x^{j}\frac{A(x)}{|A(x)|}\frac{P(x)}{Q(x)^{2}},$ (8)
$\frac{\partial F}{\partial x}=\frac{\partial P(x)}{\partial x}\frac{1}{Q(x)}-\frac{\partial Q(x)}{\partial x}\frac{P(x)}{Q^{2}(x)}.$ (9)
@@
Equations 7, 8, and 9 provides the element-wise gradients; however, since each group of size $d_{g}$, each batch of size $B$ and each element in the sequence of size $N$ share an identical set of coefficients, the contributions of each must be accumulated on $\frac{\partial F}{\partial a_{i}}$ and $\frac{\partial F}{\partial b_{j}}$. Let us denote $\frac{\partial F}{\partial a_{g,i}}$ and $\frac{\partial F}{\partial b_{g,j}}$ as the gradient for coefficient $i$ and $j$ of group $g$. Then we can express this accumulation in Equations 10 and 11:
$\frac{\partial F}{\partial a_{g,i}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial F(X_{i^{\prime},j^{\prime},g\cdot d_{g}+k^{\prime}})}{\partial a_{g,i}},$ (10)
$\frac{\partial F}{\partial b_{g,j}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial F(X_{i^{\prime},j^{\prime},g\cdot d_{g}+k^{\prime}})}{\partial b_{g.j}}.$ (11)
@@
There are multiple different methods for accumulating these gradients in a kernel, and identifying the optimal accumulation procedure is crucial for a fast training algorithm. We will now demonstrate where the standard group-wise rational function method falls short in this regard.
Analysis of KAT group-wise rational function
@@
Provided an input $\mathbf{X}\in\mathbb{R}^{B\times N\times d}$, an output gradient $\mathbf{dO}\in\mathbb{R}^{B\times N\times d}$ and coefficients $\mathbf{A}\in\mathbb{R}^{n_{g}\times m}$ and $\mathbf{B}\in\mathbb{R}^{n_{g}\times n}$, we want to compute the element-wise gradients $\frac{\partial F}{\partial\mathbf{A}}$ ($\mathbf{dA}$), $\frac{\partial F}{\partial\mathbf{B}}$ ($\mathbf{dB})$, and $\frac{\partial F}{\partial\mathbf{X}}$ ($\mathbf{dX}$). In Algorithm 1, we provide the standard method for computing each of these gradients in the KAT. The algorithm is structured sequentially, and all the on-chip computation is done in the subroutine calls, following Equations 7, 8, and 9, for readability.
Algorithm 1 The KAT group-wise rational function backward pass.
[…]
@@
3:Divide $\mathbf{X}$ and dO into $T=\lceil(B\cdot N\cdot d/S_{\text{block}}\rceil$ blocks $\mathbf{X}_{1},...,\mathbf{X}_{T}$ and $\mathbf{dO}_{1},...,\mathbf{dO}_{T}$ each of size $S_{\text{block}}$.
@@
10:  Load $\mathbf{A}_{k,1:m}$ and $\mathbf{B}_{k,1:n}$ from global memory
11:  $\mathbf{dX}_{i,j}\leftarrow\textbf{Comp\_dX}(\mathbf{A}_{k,1:m},\mathbf{B}_{k,1:n},\mathbf{X}_{i,j},\mathbf{dO}_{i,j})$
12:  $\mathbf{dA}^{\prime}_{k,1:m}\leftarrow\textbf{Comp\_dA}(\mathbf{A}_{k,1:m},\mathbf{B}_{k,1:n},\mathbf{X}_{i,j},\mathbf{dO}_{i,j})$
13:  $\mathbf{dB}^{\prime}_{k,1:n}\leftarrow\textbf{Comp\_dB}(\mathbf{A}_{k,1:m},\mathbf{B}_{k,1:n},\mathbf{X}_{i,j},\mathbf{dO}_{i,j})$
14:  Load $\mathbf{dA}_{k,1:m}$ from global memory, set $\mathbf{dA}_{k,1:m}\leftarrow\mathbf{dA}_{k,1:m}+\mathbf{dA}^{\prime}_{k,1:m}$, write back to global memory
15:  Load $\mathbf{dB}_{k,1:n}$ from global memory, set $\mathbf{dB}_{k}\leftarrow\mathbf{dB}_{k}+\mathbf{dB}^{\prime}_{k}$, write back to global memory
@@
At a high level, Algorithm 1 iterates over the $T$ blocks in the grid, where each block has $S_{\text{block}}$ threads that compute the associated gradients for $\mathbf{dA}$, $\mathbf{dB}$, $\mathbf{dX}$. In the case of $\mathbf{dA}$ and $\mathbf{dB}$, for each of these computations, the thread performs an atomic add, which requires reading from global memory, accumulating the value, and then storing the result back to global memory as an indivisible unit of work. However, since each of these threads must perform an atomic add for each of the $m+n$ coefficients, the atomic adds must occur sequentially. As such, for each of these atomic adds, a stall occurs. Since multiple threads within the same block and across different blocks may be writing to the same location, there will be extreme resource contention. Such resource contention inevitably leads to high latency and stalls without utilizing the available bandwidth, as we demonstrated in Section 3.
@@
Let us analyze the total number of global memory accesses incurred by Algorithm 1. We can see that each $\mathbf{X}_{i}$ and $\mathbf{dO}_{i}$ requires $S_{\text{block}}$ reads from global memory. Then writing $\mathbf{dX}_{i}$ back to global memory requires $S_{\text{block}}$ writes to global memory. Since these 3 memory actions occur $T$ times they require $3TS_{\text{block}}=3\frac{B\cdot N\cdot d}{S_{\text{block}}}\cdot S_{\text{block}}=3\cdot B\cdot N\cdot d$ global memory accesses. Regarding the coefficients, each element in the output $dX$ requires loading $m+n$ coefficients from global memory. It also requires the aforementioned atomic add, which includes an additional $(m+n)$ reads and $(m+n)$ writes to global memory. Thus, since there are $T$ blocks of $S_{\text{block}}$ threads, there are $TS_{\text{block}}\cdot 3(m+n)=3(m+n)\cdot B\cdot N\cdot d$ accesses. As a result, there are $3(m+n+1)\cdot B\cdot N\cdot d$ global memory accesses in total. To reiterate, the memory accesses for the coefficients for the atomic add are much worse due to the additional contention with other threads.
FlashKAT group-wise rational function
@@
We now present an optimized method for computing the backward pass of the group-wise rational function in Algorithm 2. To reduce global memory accesses, we restructure the grid from 1D to 2D. Such a restructuring mirrors the format of Equation 10, allowing an easier mapping of it to the algorithm. The first dimension of our reconstructed grid is allocated for resolving the $B$ and $N$ summations, whereas the second dimension is responsible for the $d_{g}$ summation. The size of this first dimension in the grid is $B\cdot N/S_{\text{block}}$, whereas the size of the second dimension is $n_{g}$ (the total number of groups in GR-KAN).
@@
Algorithm 2 The FlashKAT group-wise rational function backward pass.
[…]
2:Set block sizes to $S_{\text{block}}$ and $d_{g}=d/n_{g}$.
@@
3:Divide $\mathbf{X}$ and $\mathbf{dO}$ into $\lceil(B\cdot N\cdot d/(S_{\text{block}}\cdot d_{g})\rceil$ blocks $\mathbf{X}_{i,j}\in\mathbb{R}^{S_{\text{block}}\times d_{g}}$ and $\mathbf{dO}_{i,j}\in\mathbb{R}^{S_{\text{block}}\times d_{g}}$ of a $T\times n_{g}$ grid where $T=\lceil(B\cdot N/S_{\text{block}}\rceil$.
@@
9:  Load $\mathbf{X}_{i,j}$ from global memory
@@
14:   $\mathbf{dX}_{i,j,k,l}\leftarrow\textbf{Comp\_dX}(\mathbf{A}_{j,1:m},\mathbf{B}_{j,1:n},\mathbf{X}_{i,j,k,l},\mathbf{dO}_{i,j,k,l})$
15:   $\mathbf{dA}^{\prime}_{j}\leftarrow\mathbf{dA}^{\prime}_{j}+\textbf{Comp\_dA}(\mathbf{A}_{j,1:m},\mathbf{B}_{j,1:n},\mathbf{X}_{i,k},\mathbf{dO}_{i,j,k,l})$
16:   $\mathbf{dB}^{\prime}_{j}\leftarrow\mathbf{dB}^{\prime}_{j}+\textbf{Comp\_dB}(\mathbf{A}_{j,1:m},\mathbf{B}_{j,1:n},\mathbf{X}_{i,j,k,l},\mathbf{dO}_{i,j,k,l})$
@@
19:  Load $\mathbf{dA}_{j}$ from global memory, set $\mathbf{dA}_{j}\leftarrow\mathbf{dA}_{j}+\mathbf{dA}^{\prime}_{j}$, write back to global memory
20:  Load $\mathbf{dB}_{j}$ from global memory, set $\mathbf{dB}_{j}\leftarrow\mathbf{dB}_{j}+\mathbf{dB}^{\prime}_{j}$, write back to global memory
@@
Algorithm 2 does not change the memory accesses for reads and writes for $\mathbf{dX},\mathbf{X}$ and $\mathbf{dO}$; however, it heavily influences them for $\mathbf{A}$, $\mathbf{B}$, $\mathbf{dA}$, and $\mathbf{dB}$. At a high level, we can see that Algorithm 2 iterates over the grid of $T\cdot n_{g}$ blocks. For each of these blocks, it loads a single group of coefficients for $\mathbf{A}_{j}$ and $\mathbf{B}_{j}$, which it then reuses for all computations in the block. As a result, rather than each thread performing an atomic add for each value in $\mathbf{dA}_{j}$ and $\mathbf{dB}_{j}$, the algorithm can instead accumulate all the contributions for $\mathbf{dA}_{j}$ and $\mathbf{dB}_{j}$ in the block before doing an atomic add. As a result, Algorithm 2 reduces the total amount of memory contention and global memory accesses.
We now analyze the total global memory accesses for Algorithm 2. Each block first loads one set of coefficients from the global memory associated with the group to which the block is assigned. Since there is only 1 set of coefficients per block, this action only requires $m+n$ accesses. Then the algorithm loads the associated $\mathbf{X}_{i,j}$ and $\mathbf{dO}_{i,j}$ from global memory, each of which is of size $S_{\text{block}}\cdot d_{g}$. After computing $\mathbf{dX}_{i,j}$, its $S_{\text{block}}\cdot d_{g}$ values are written back to global memory. Then, an atomic add loads $m+n$ values from $\mathbf{dA}_{j}$ and $\mathbf{dB_{j}}$, accumulates them with $\mathbf{dA}^{\prime}_{j}$ and $\mathbf{dB}^{\prime}_{j}$, and writes them back to global memory. Since all these operations occur across the $T\cdot n_{g}$ blocks, there are $T\cdot n_{g}(3S_{\text{block}}\cdot d_{g}+3(m+n))=3\frac{B\cdot N}{S_{\text{block}}}\cdot\frac{d}{d_{g}}(S_{\text{block}}\cdot d_{g}+(m+n))=3(\frac{m+n}{S_{\text{block}}\cdot d_{g}}+1)\cdot B\cdot N\cdot d$ global memory access. Compared to Algorithm 1, we can see that Algorithm 2 reduces the number of global memory accesses and subsequently the number of atomic adds by a factor of $\frac{1}{S_{\text{block}}\cdot d_{g}}$.